OpenStudy (anonymous):

NEED HELP FACTORING: (x+8)^2 - 3

OpenStudy (confluxepic):

Ok.

OpenStudy (confluxepic):

I can try.

OpenStudy (mathstudent55):

Is the problem copied correctly, or is there anything missing?

OpenStudy (studygurl14):

\[(a+b)^2 = a^2+2ab+b^2\] \[(x+8)^2\] a = x, b = 8 \[a^2 + 2ab+b^2\] \[(x)^2+2(x)(8)+(8)^2\] \[x^2+16x+64\] Then minus 3 \[x^2+16x+64 - 3 = x^2+16x+61\]

OpenStudy (anonymous):

well it says a to factor y=(x+8)^2 - 3

OpenStudy (mathstudent55):

OK. This is how I'd do it. (x + 8)^2 is the square of x + 8. 3 is the square of \(\sqrt{3} \) That means you can rewrite this as the difference of two squares: \((x + 8)^2 - (\sqrt{3})^2\)

OpenStudy (mathstudent55):

The difference of two squares factors like this: \(a^2 - b^2 = (a + b)(a - b)\) Now apply this to your problem.

OpenStudy (mathstudent55):

\((x + 8)^2 - (\sqrt{3})^2 = (x + 8 + \sqrt{3})(x + 8 - \sqrt{3} )\)